# thanks: M.C. and D.P. contributed equally to this work

### v2

@@ Abstract
It is a common belief that a theory of quantum gravity should ultimately cure curvature singularities which are inevitable within General Relativity, and plague for instance the Schwarzschild and Kerr metrics, usually considered as prototypes for primordial black holes (PBHs) as dark matter (DM) candidates. We continue our study, initiated in a companion paper, of non-singular objects as PBHs, considering three regular non-tr-symmetric metrics, all of which are one-parameter extensions of the Schwarzschild space-time: the Simpson-Visser, Peltola-Kunstatter, and D’Ambrosio-Rovelli space-times, with the latter two motivated by loop quantum gravity. We study evaporation constraints on PBHs described by these regular metrics, deriving upper limits on $f_{\text{pbh}}$, the fraction of DM in the form of PBHs. Compared to their Schwarzschild counterparts, these limits are weaker, and result in a larger asteroid mass window where all the DM can be in the form of PBHs. Our work demonstrates as a proof-of-principle that quantum gravity-inspired space-times can simultaneously play an important role in the resolution of singularities and in the DM problem.
@@ I Introduction
The collapse of large density perturbations upon horizon re-entry in the early Universe can lead to the formation of hypothetical relics known as primordial BHs (PBHs), whose role as potential DM candidates has long been recognized Chapline (1975); Meszaros (1975); Khlopov and Polnarev (1980); Khlopov et al. (1985); Ivanov et al. (1994); Choudhury and Mazumdar (2014); Belotsky et al. (2014); Bird et al. (2016); Clesse and García-Bellido (2017); Poulin et al. (2017); Raccanelli et al. (2018); Bernal et al. (2017); Clesse and García-Bellido (2018); Kohri and Terada (2018); Murgia et al. (2019); Carr et al. (2021a); Serpico et al. (2020); De Luca et al. (2020a, b, c); Carr et al. (2021b); Bhagwat et al. (2021); De Luca et al. (2021a); Wong et al. (2021); Carr et al. (2021c); Domènech et al. (2021); De Luca et al. (2021b); Arbey et al. (2021a); Franciolini et al. (2022a); De Luca et al. (2021c); Cheek et al. (2022a, b); Heydari and Karami (2022a); Dvali et al. (2021); Heydari and Karami (2022b); De Luca et al. (2022); Bhaumik et al. (2022); Cai et al. (2022); Oguri et al. (2023); Franciolini et al. (2022b); Mazde and Visinelli (2023); Cai et al. (2023); Fu and Wang (2023); Choudhury et al. (2024a); Papanikolaou (2023); Choudhury et al. (2023a, b); Musco et al. (2024); Yuan et al. (2024); Choudhury et al. (2023c); Ghoshal et al. (2023); Cai et al. (2024a); Choudhury et al. (2024b); Huang et al. (2024a); Choudhury et al. (2023d); Bhattacharya et al. (2024); Heydari and Karami (2024a, b); Choudhury et al. (2024c, 2023e); Ghoshal and Strumia (2024); Huang et al. (2024b); Huang and Piao (2024); Choudhury et al. (2024d); Thoss et al. (2024); Papanikolaou et al. (2024a); Choudhury (2024); Choudhury et al. (2024e); Papanikolaou et al. (2024b); Yin and Visinelli (2024); Choudhury et al. (2024f); Heydari and Karami (2024c); Dvali et al. (2024); Boccia et al. (2024); Huang et al. (2024c); Choudhury et al. (2024g); Yang et al. (2024) (for recent reviews on the topic, see Refs. Khlopov (2010); Carr et al. (2016a); Green and Kavanagh (2021); Carr and Kuhnel (2020); Villanueva-Domingo et al. (2021); Carr and Kuhnel (2022); Bird et al. (2023); Carr et al. (2024); Arbey (2024); Choudhury and Sami (2024)). A wide range of observations (mainly of astrophysical nature) severely limit the ability of PBHs to account for the entire DM component: in practice, this is potentially possible (although this possibility is not completely free of debates) only in the so-called “asteroid mass window”, i.e. $10^{14}\,{\text{kg}}\lesssim M_{\text{pbh}}\lesssim 10^{20}\,{\text{kg}}$, with lighter and heavier PBHs being tightly constrained by observational signatures of their evaporation and microlensing respectively Katz et al. (2018); Smyth et al. (2020); Coogan et al. (2021); Ray et al. (2021); Auffinger (2022); Ghosh and Mishra (2024); Miller et al. (2022); Branco et al. (2023); Bertrand et al. (2023); Tran et al. (2023); Gorton and Green (2024); Dent et al. (2024); Tamta et al. (2024); Tinyakov (2024); Loeb (2024). However, it is important to note that virtually all constraints on PBHs, including those determining the existence and extension of the asteroid mass window, are subject to the underlying assumption about these object being either Schwarzschild or Kerr BHs Khlopov (2010); Carr et al. (2016a); Green and Kavanagh (2021); Carr and Kuhnel (2020); Villanueva-Domingo et al. (2021); Carr and Kuhnel (2022); Bird et al. (2023); Carr et al. (2024); Arbey (2024); Choudhury and Sami (2024). This assumption is perfectly reasonable from the phenomenological and observational point of view, but at the same time may be cause of some apprehensiveness on the more theoretical side. In fact, these metrics feature pathological curvature singularities, whose existence is virtually inevitable in General Relativity (GR), and is at the essence of the well-known singularity problem Ansoldi (2008); Nicolini (2009); Sebastiani and Zerbini (2022); Torres (2022); Lan et al. (2023).
@@
To be concrete, in what follows we will consider three regular, static spherically symmetric space-times, characterized by an additional regularizing parameter $\ell$, and recovering the Schwarzschild space-time in the $\ell\to 0$ limit. All three space-times enjoy quite different properties compared to the phenomenological ones considered in our companion paper Calzà et al. (2024a). The first metric we consider is the so-called Simpson-Visser metric: this is arguably one of the best known black-bounce space-times, and interpolates between the Schwarzschild metric, regular BHs, and traversable wormholes. The other two metrics are instead deeply rooted within Loop Quantum Gravity (LQG) Ashtekar (1986); Rovelli (1998); Ashtekar and Bianchi (2021): arguably one of the leading QG approaches, LQG is a fully non-perturbative and manifestly background-independent approach towards a consistent theory of QG, wherein space-time is fundamentally discrete (see e.g. Refs. Bojowald (2001, 2005); Modesto (2008); Engle et al. (2008); Modesto (2009); Bianchi et al. (2011); Singh and Vidotto (2011); Bojowald and Paily (2012); Cailleteau et al. (2012); Agullo et al. (2013); Haggard and Rovelli (2015); Odintsov and Oikonomou (2014); Odintsov et al. (2015); Haro et al. (2015); Odintsov and Oikonomou (2016); De Haro et al. (2018); de Haro et al. (2019); Odintsov and Oikonomou (2018); Benetti et al. (2019); Casalino et al. (2020); Bouhmadi-López et al. (2020); Graef et al. (2020); Brahma et al. (2021); Barboza et al. (2022); S. Vicente et al. (2022); Afrin et al. (2023); Yan et al. (2023); Kumar et al. (2023); Jiang et al. (2024a, b) for various follow-up studies and applications). More specifically, the two regular LQG-motivated metrics we analyze as candidates for DM in the form of PRBHs are the Peltola-Kunstatter Peltola and Kunstatter (2009a, b) and D’Ambrosio-Rovelli space-times Bianchi et al. (2018); D’Ambrosio and Rovelli (2018). As a cautionary note, we remark that ours is to be intended as a pilot study in this direction, and that much more follow-up work is needed before primordial regular BHs as DM candidates are characterized to the same extent as their Schwarzschild counterparts.  [^1] As far as we are aware, only four earlier works considered primordial regular BHs as DM Dymnikova and Khlopov (2015); Pacheco (2018); Arbey and Auffinger (2021); Arbey et al. (2022); Banerjee et al. (2024), but mostly focused on aspects other than the ones studied here (with the exception of Refs. Arbey and Auffinger (2021); Arbey et al. (2022), which however studied a different LQG-inspired metric, albeit reaching conclusions qualitatively similar to ours).
@@ II Regular black holes
$T=\frac{f^{\prime}(r)}{4\pi\sqrt{-g(r)f(r)}}|_{r_{H}}\,,$ (5)
@@ II.1 Simpson-Visser space-time
The SV space-time encompasses a rich phenomenology, as it interpolates between the Schwarzschild BH ($\ell=0$), a regular BH with a one-way space-like throat ($20<\ell/M<$ 2), a one-way WH with an extremal null throat ($\ell/M=2$), and a traversable WH with a two-way time-like throat ($\ell/M>2$). This metric has been the subject of several follow-up studies (see e.g. Refs. Tsukamoto (2021); Mazza et al. (2021); Islam et al. (2021); Guerrero et al. (2021); Bambhaniya et al. (2022); Yang et al. (2022); Riaz et al. (2022); Arora et al. (2024); Jha (2023); Jha and Rahaman (2023)) and, while originally introduced on phenomenological grounds, can potentially originate as a solution of GR coupled to non-linear electrodynamics in the presence of a minimally coupled phantom scalar field Bronnikov and Walia (2022).
@@ III.1 Greybody factors
To compute the GBFs, we make use of the shooting method, widely used earlier in similar contexts (see e.g. Refs. Rosa (2017, 2013); Calzà et al. (2021); Calzá (2023); Calzà and Rosa (2022); Calzà et al. (2023); Calzà and Rosa (2023a, b)), including in our companion paper Calzà et al. (2024a). We begin by defining the rescaled coordinate $x$:
@@ III.2 Evaporation spectra
Figure 2: Primary photon spectra resulting from the evaporation of Simpson-Visser black holes of mass $10^{13}\,{\text{kg}}$ for different values of the regularizing parameter $\ell$ (normalized by the horizon radius $r_{H}$): $\ell/r_{H}=0.3$ (red dotted curve), $0.6$ (green dashed curve), and $0.9$ (magenta dash-dotted curve). The blue solid curve corresponds to the case $\ell/r_{H}=0$, which recovers the Schwarzschild black hole.
@@ III.3 Evaporation constraints
In the mass range $10^{10}\,{\text{kg}}\lesssim M_{\text{pbh}}\lesssim 10^{15}\,{\text{kg}}$, the dominant constraints on the PBH abundance come from measurements of the extragalactic photon background Auffinger (2023), and more precisely of the diffuse extragalactic $\gamma$-ray background (EGRB) in the energy range $100\,{\text{keV}}\lesssim E_{\gamma}\lesssim 5\,{\text{GeV}}$, which can be directly compared against theoretical expectations for the PBH Hawking evaporation spectra. Of particular interest to us is the fact that the lower edge of the asteroid mass window, where PBHs could make up the entire DM, is precisely set by evaporation constraints. In what follows, we set evaporation constraints on the fraction of DM in the form of PBHs $f_{\text{pbh}}(M)\equiv\Omega_{\text{pbh}}/\Omega_{\text{dm}}$, assuming that PRBHs are described by the three metrics discussed so far. We work under the same set of approximations adopted in our companion paper Calzà et al. (2024a). Namely, we assume that PRBHs are isotropically distributed on sufficiently large scales and cluster in the galactic halo in the same way as other forms as DM, we only compute the primary photon spectrum, and finally we assume a monochromatic mass distribution (see e.g. Refs. Kühnel et al. (2016); Kühnel and Freese (2017); Carr et al. (2017); Raidal et al. (2017); Bellomo et al. (2018); Lehmann et al. (2018); Carr and Kuhnel (2019); Gow et al. (2020); De Luca et al. (2020d); Gow et al. (2022); Ashoorioon et al. (2021); Bagui and Clesse (2022); Mukhopadhyay et al. (2022); Papanikolaou (2022); Cai et al. (2024b) for studies on the effects of an extended mass distribution). We refer the reader to Sec. IIIC of our companion paper Calzà et al. (2024a) for a more detailed discussion of why these, which are clearly all approximations, are appropriate for the scope of our work (while allowing for a more direct comparison to earlier works, including our companion paper).
@@
with $\rho_{\text{crit},0}=3H_{0}^{2}/8\pi G$ being the present-day critical density, $H_{0}$ the Hubble constant, and $\Omega_{\text{dm}}$ the present-day DM density parameter. In what follows, in order to specify $H(z)$, $\rho_{\text{crit}}$, and $\Omega_{\text{dm}}$, we adopt the same spatially flat $\Lambda$CDM cosmological model used by the seminal Ref. Carr et al. (2010). This allows us to have a reliable reference against which we can cross-check our limits on $f_{\text{pbh}}$ in the Schwarzschild case ($\ell\to 0$), although we stress that the choice of underlying cosmology does not play a significant role in determining our constraints.

### v3

@@ Abstract
It is a common belief that a theory of quantum gravity should ultimately cure curvature singularities which are inevitable within General Relativity, and plague for instance the Schwarzschild and Kerr metrics, usually considered as prototypes for primordial black holes (PBHs) as dark matter (DM) candidates. We continue our study, initiated in a companion paper, of non-singular objects as PBHs, considering three regular non-tr (non-time-radial)-symmetric metrics, all of which are one-parameter extensions of the Schwarzschild space-time: the Simpson-Visser, Peltola-Kunstatter, and D’Ambrosio-Rovelli space-times, with the latter two motivated by loop quantum gravity. We study evaporation constraints on PBHs described by these regular metrics, deriving upper limits on $f_{\text{pbh}}$, the fraction of DM in the form of PBHs. Compared to their Schwarzschild counterparts, these limits are weaker, and result in a larger asteroid mass window where all the DM can be in the form of PBHs, with the lower edge moving potentially more than an order of magnitude. Our work demonstrates as a proof-of-principle that quantum gravity-inspired space-times can simultaneously play an important role in the resolution of singularities and in the DM problem.
@@ I Introduction
The collapse of large density perturbations upon horizon re-entry in the early Universe can lead to the formation of hypothetical relics known as primordial BHs (PBHs), whose role as potential DM candidates has long been recognized Chapline (1975); Meszaros (1975); Khlopov and Polnarev (1980); Khlopov et al. (1985); Ivanov et al. (1994); Choudhury and Mazumdar (2014); Belotsky et al. (2014); Bird et al. (2016); Clesse and García-Bellido (2017); Poulin et al. (2017); Raccanelli et al. (2018); Bernal et al. (2017); Clesse and García-Bellido (2018); Kohri and Terada (2018); Liu et al. (2019a, b); Murgia et al. (2019); Carr et al. (2021a); Liu et al. (2020); Hertzberg et al. (2020); Serpico et al. (2020); De Luca et al. (2020a, b, c); Carr et al. (2021b); Bhagwat et al. (2021); De Luca et al. (2021a); Wong et al. (2021); Carr et al. (2021c); Domènech et al. (2021); De Luca et al. (2021b); Arbey et al. (2021a); Franciolini et al. (2022a); De Luca et al. (2021c); Cheek et al. (2022a, b); Heydari and Karami (2022a); Dvali et al. (2021); Heydari and Karami (2022b); De Luca et al. (2022); Liu et al. (2023a); Saha and Laha (2022); Bhaumik et al. (2022); Anchordoqui et al. (2022); Cai et al. (2022); Oguri et al. (2023); Franciolini et al. (2022b); Mazde and Visinelli (2023); Cai et al. (2023); Anchordoqui et al. (2023); Liu et al. (2023b); Fu and Wang (2023); Choudhury et al. (2024a); Papanikolaou (2023); de Freitas Pacheco et al. (2023); Choudhury et al. (2023a, b); Musco et al. (2024); Yuan et al. (2024); Choudhury et al. (2023c); Ghoshal et al. (2023); Cai et al. (2024a); Choudhury et al. (2024b); Huang et al. (2024a); Choudhury et al. (2024c); Bhattacharya et al. (2024); Heydari and Karami (2024a, b); Choudhury et al. (2024d, 2023d); Ghoshal and Strumia (2024); Huang et al. (2024b); Huang and Piao (2024); Anchordoqui et al. (2024a); Choudhury et al. (2024e); Thoss et al. (2024); Papanikolaou et al. (2024a); Choudhury (2024); Choudhury et al. (2024f); Anchordoqui et al. (2024b); Papanikolaou et al. (2024b); Yin and Visinelli (2024); Choudhury et al. (2024g); Heydari and Karami (2024c); Dvali et al. (2024); Boccia et al. (2024); Huang et al. (2024c); Choudhury et al. (2024h); Anchordoqui et al. (2024c); Yang et al. (2024); Saha et al. (2024); Anchordoqui et al. (2024d); Chen et al. (2024); Dai and Stojkovic (2024); Huang et al. (2024d); Zantedeschi and Visinelli (2024); Chianese et al. (2024); Barker et al. (2024); Borah and Das (2024); Huang et al. (2024e) (for recent reviews on the topic, see Refs. Khlopov (2010); Carr et al. (2016a); Green and Kavanagh (2021); Carr and Kuhnel (2020); Villanueva-Domingo et al. (2021); Carr and Kuhnel (2022); Bird et al. (2023); Carr et al. (2024); Arbey (2024); Choudhury and Sami (2024)). A wide range of observations (mainly of astrophysical nature) severely limit the ability of PBHs to account for the entire DM component: in practice, this is potentially possible (although this possibility is not completely free of debates) only in the so-called “asteroid mass window”, i.e. $10^{17}\,{\text{g}}\lesssim M_{\text{pbh}}\lesssim 10^{23}\,{\text{g}}$, with lighter and heavier PBHs being tightly constrained by observational signatures of their evaporation and microlensing respectively Katz et al. (2018); Bai and Orlofsky (2019); Smyth et al. (2020); Coogan et al. (2021); Ray et al. (2021); Auffinger (2022); Ghosh and Mishra (2024); Miller et al. (2022); Branco et al. (2023); Bertrand et al. (2023); Tran et al. (2024); Gorton and Green (2024); Dent et al. (2024); Tamta et al. (2024); Tinyakov (2024); Loeb (2024). However, it is important to note that virtually all constraints on PBHs, including those determining the existence and extension of the asteroid mass window, are subject to the underlying assumption about these object being either Schwarzschild or Kerr BHs Khlopov (2010); Carr et al. (2016a); Green and Kavanagh (2021); Carr and Kuhnel (2020); Villanueva-Domingo et al. (2021); Carr and Kuhnel (2022); Bird et al. (2023); Carr et al. (2024); Arbey (2024); Choudhury and Sami (2024). This assumption is perfectly reasonable from the phenomenological and observational point of view, but at the same time may be cause of some apprehensiveness on the more theoretical side. In fact, these metrics feature pathological curvature singularities, whose existence is virtually inevitable in General Relativity (GR), and is at the essence of the well-known singularity problem Ansoldi (2008); Nicolini (2009); Sebastiani and Zerbini (2022); Torres (2022); Lan et al. (2023).
@@
To be concrete, in what follows we will consider three regular, static spherically symmetric space-times, characterized by an additional regularizing parameter $\ell$, and recovering the Schwarzschild space-time in the $\ell\to 0$ limit. All three space-times enjoy quite different properties compared to the phenomenological ones considered in our companion paper Calzà et al. (2024a). The first metric we consider is the so-called Simpson-Visser metric: this is arguably one of the best known black-bounce space-times, and interpolates between the Schwarzschild metric, regular BHs, and traversable wormholes.  [^1] A traversable wormhole is one for which a particle can enter through one side of the wormhole and exit through the other. In principle, these could be traversed by a human traveler without fatal effects Morris and Thorne (1988), if certain conditions on the metric elements $g_{tt}$ and $g_{rr}$ are satisfied. In practice, whether this trip would actually be free from fatal effects depends on the question of how a human body would interact with the exotic matter which is required to sustain the wormhole, question which is presently open. The other two metrics are instead deeply rooted within Loop Quantum Gravity (LQG) Ashtekar (1986); Rovelli (1998); Ashtekar and Bianchi (2021): arguably one of the leading QG approaches, LQG is a fully non-perturbative and manifestly background-independent approach towards a consistent theory of QG, wherein space-time is fundamentally discrete (see e.g. Refs. Bojowald (2001, 2005); Modesto (2008); Engle et al. (2008); Modesto (2009); Bianchi et al. (2011); Singh and Vidotto (2011); Bojowald and Paily (2012); Cailleteau et al. (2012); Agullo et al. (2013); Haggard and Rovelli (2015); Odintsov and Oikonomou (2014); Odintsov et al. (2015); Haro et al. (2015); Odintsov and Oikonomou (2016); De Haro et al. (2018); de Haro et al. (2019); Odintsov and Oikonomou (2018); Benetti et al. (2019); Casalino et al. (2020); Bouhmadi-López et al. (2020); Graef et al. (2020); Brahma et al. (2021); Barboza et al. (2022); S. Vicente et al. (2022); Afrin et al. (2023); Yan et al. (2023); Kumar et al. (2023); Jiang et al. (2024a, b) for various follow-up studies and applications). More specifically, the two regular LQG-motivated metrics we analyze as candidates for DM in the form of PRBHs are the Peltola-Kunstatter Peltola and Kunstatter (2009a, b) and D’Ambrosio-Rovelli space-times Bianchi et al. (2018); D’Ambrosio and Rovelli (2018). As a cautionary note, we remark that ours is to be intended as a pilot study in this direction, and that much more follow-up work is needed before primordial regular BHs as DM candidates are characterized to the same extent as their Schwarzschild counterparts.  [^2] As far as we are aware, only seven earlier works considered primordial regular BHs as DM Easson (2003); Dymnikova and Khlopov (2015); Pacheco (2018); Arbey and Auffinger (2021); Arbey et al. (2022); Banerjee et al. (2024); Davies et al. (2024), but mostly focused on aspects other than the ones studied here (with the exception of Refs. Arbey and Auffinger (2021); Arbey et al. (2022), which however studied a different LQG-inspired metric, albeit reaching conclusions qualitatively similar to ours).
@@ II Regular black holes
$T=\sqrt{\frac{g(r)}{f(r)}}\frac{f^{\prime}(r)}{4\pi}|_{r_{H}}\,,$ (5)
@@ II.1 Simpson-Visser space-time
The SV space-time encompasses a rich phenomenology, as it interpolates between the Schwarzschild BH ($\ell=0$), a regular BH with a one-way space-like throat ($20<\ell/M<$ 2), a one-way WH with an extremal null throat ($\ell/M=2$), and a traversable WH with a two-way time-like throat ($\ell/M>2$).  [^5] These conditions are derived by looking at a) the existence (or not) of horizons from the zeros of $g^{rr}$, which implies $r_{H}=\sqrt{4M^{2}-\ell^{2}}$, and therefore $\ell<2M$ in order to have a BH, and b) the sign of the “coordinate speed of light” $|dr/dt|=1-2M/\sqrt{r^{2}+\ell^{2}}$, which determines whether the wormhole for $\ell\geq 2M$ is traversable ($dr/dt\neq 0$ for all $r$) or not ($dr/dt\to 0$ as $r\to 0$) – for a more detailed explanation, see the discussion above Eq. (2.6) of Ref. Simpson and Visser (2019). This metric has been the subject of several follow-up studies (see e.g. Refs. Tsukamoto (2021); Mazza et al. (2021); Shaikh et al. (2021); Islam et al. (2021); Guerrero et al. (2021); Bambhaniya et al. (2022); Yang et al. (2022); Riaz et al. (2022); Arora et al. (2024); Jha (2023); Jha and Rahaman (2023)) and, while originally introduced on phenomenological grounds, can potentially originate as a solution of GR coupled to non-linear electrodynamics in the presence of a minimally coupled phantom scalar field Bronnikov and Walia (2022).
@@ III.1 Greybody factors
To compute the GBFs, we make use of the shooting method, widely used earlier in similar contexts (see e.g. Refs. Rosa (2017, 2013); Calzà et al. (2021); Calzá (2023); Calzà and Rosa (2022); Calzà et al. (2024c); Calzà and Rosa (2024, 2023)), including in our companion paper Calzà et al. (2024a). We begin by defining the rescaled coordinate $x$:
@@
Figure 2: Greybody factors $\Gamma_{l=1}^{s=1}$ as a function of $\omega/M$ for Schwarzschild BHs (black curve), as well as the Simpson-Visser (blue solid curve), Peltola-Kunstatter (red dashed curve), and D’Ambrosio-Rovelli (green dotted curve) regular space-times. For illustrative purposes we only plot $\Gamma_{l=1}^{s=1}$, since we are interested in photons ($s=1$) and the dominant emission mode is the $l=1$ one. We have fixed the regularizing parameter to $\ell=0.3r_{H}$ for all three regular space-times. We see that in all three cases the GBFs are consistently (slightly) higher than their Schwarzschild counterparts.
In Fig. 2 we plot the $\Gamma_{l=1}^{s=1}$ GBFs for the Schwarzschild BH and the three regular space-times we study, focusing for illustrative purposes on the $\Gamma_{l=1}^{s=1}$ GBF and fixing $\ell=0.3r_{H}$ for all three regular space-times. We observe that the GBFs are slightly higher than their Schwarzschild counterparts (by $\lesssim 10\%$ at most), and asymptote to the latter for both $\omega/M\lesssim 0.3$ and $\omega/M\gtrsim 0.7$.
@@ III.2 Evaporation spectra
Figure 3: Primary photon spectra resulting from the evaporation of Simpson-Visser black holes of mass $10^{16}\,{\text{g}}$ for different values of the regularizing parameter $\ell$ (normalized by the horizon radius $r_{H}$): $\ell/r_{H}=0.3$ (red dotted curve), $0.6$ (green dashed curve), and $0.9$ (magenta dash-dotted curve). The blue solid curve corresponds to the case $\ell/r_{H}=0$, which recovers the Schwarzschild black hole.
@@
Examples of the resulting evaporation spectra are provided in Figs. 34, and 5 for a representative PBH of mass $M_{\text{pbh}}=10^{16}\,{\text{g}}$, located somewhat halfway in the mass range of interest (although the features we discuss shortly do not depend on the chosen mass). For the SV and DR PRBHs, increasing $\ell$ leads to the intensity of the spectra decreasing at all energies. This feature confirms the intuition raised when we discussed the temperatures of these BHs (see Fig. 1), which we observed to decrease with increasing $\ell$. However, we remark that a decrease in $T$ alone is not sufficient to draw this conclusion, since the GBFs also enter in Eq. (34). Nevertheless, one can expect the temperature to play a more important role, as it enters exponentially into Eq. (34), unlike the GBFs which enter linearly.
@@ III.3 Evaporation constraints
In the mass range $10^{13}\,{\text{g}}\lesssim M_{\text{pbh}}\lesssim 10^{18}\,{\text{g}}$, the dominant constraints on the PBH abundance come from measurements of the extragalactic photon background Auffinger (2023), and more precisely of the diffuse extragalactic $\gamma$-ray background (EGRB) in the energy range $100\,{\text{keV}}\lesssim E_{\gamma}\lesssim 5\,{\text{GeV}}$, which can be directly compared against theoretical expectations for the PBH Hawking evaporation spectra. Of particular interest to us is the fact that the lower edge of the asteroid mass window, where PBHs could make up the entire DM, is precisely set by evaporation constraints. In what follows, we set evaporation constraints on the fraction of DM in the form of PBHs $f_{\text{pbh}}(M)\equiv\Omega_{\text{pbh}}/\Omega_{\text{dm}}$, assuming that PRBHs are described by the three metrics discussed so far. We work under the same set of approximations adopted in our companion paper Calzà et al. (2024a). Namely, we assume that PRBHs are isotropically distributed on sufficiently large scales and cluster in the galactic halo in the same way as other forms as DM, we only compute the primary photon spectrum, and finally we assume a monochromatic mass distribution (see e.g. Refs. Kühnel et al. (2016); Kühnel and Freese (2017); Carr et al. (2017); Raidal et al. (2017); Bellomo et al. (2018); Lehmann et al. (2018); Carr and Kuhnel (2019); Gow et al. (2020); De Luca et al. (2020d); Gow et al. (2022); Ashoorioon et al. (2021); Bagui and Clesse (2022); Mukhopadhyay et al. (2022); Papanikolaou (2022); Cai et al. (2024b) for studies on the effects of an extended mass distribution). We refer the reader to Sec. IIIC of our companion paper Calzà et al. (2024a) for a more detailed discussion of why these, which are clearly all approximations, are appropriate for the scope of our work (while allowing for a more direct comparison to earlier works, including our companion paper).
@@
with $\rho_{\text{crit},0}=3H_{0}^{2}/8\pi G$ being the present-day critical density, $H_{0}$ the Hubble constant, and $\Omega_{\text{dm}}$ the present-day DM density parameter. In what follows, in order to specify $H(z)$, $\rho_{\text{crit}}$, and $\Omega_{\text{dm}}$, we adopt the same spatially flat $\Lambda$CDM cosmological model used by the seminal Ref. Carr et al. (2010). This allows us to have a reliable reference against which we can cross-check our limits on $f_{\text{pbh}}$ in the Schwarzschild case ($\ell\to 0$), although we stress that the choice of underlying cosmology does not play a significant role in determining our constraints. Analogously to our companion paper, we focus on the mass range $M_{\text{pbh}}>10^{15}\,{\text{g}}$. A Schwarzschild PBH of this mass has a lifetime which is much longer than the age of the Universe, and is therefore far from having fully evaporated today. Importantly, it has only lost a negligible fraction of its mass from formation until today, and it is therefore safe to approximate these PBHs as being quasi-static throughout the lifetime of the Universe, while denoting by $M_{\text{pbh}}$ the values of the PBH mass both at formation and today. These considerations hold even more strongly for the regular BHs we consider, as these are colder and hence longer lived compared to their Schwarzschild counterparts at a given mass. We refer the reader to Appendix B of our companion paper Calzà et al. (2024a) for further discussions on this point.